OpenStudy (cookie_2046):

What is the value of 54 – 4a2 + 3b3 when a = −2 and b = 4?

OpenStudy (anonymous):

just sub in your values 54 - 4(-2^2) + 3(4^3)....can you do the math now ?

OpenStudy (anonymous):

yeah just what kelliegirl13 said

OpenStudy (cookie_2046):

The answer is 230 right?

OpenStudy (cookie_2046):

@kelliegirl33

OpenStudy (anonymous):

let me check.. 54 - 4(-2^2) + 3(4^3) = 54 - 4(4) + 3(64) = 54 - 16 + 192 = 246 - 16 = 230......yep...thats the same thing I got :)

OpenStudy (cookie_2046):

Alright well thank you

OpenStudy (anonymous):

sure thing :)